# UniFoil: A Universal Dataset of Airfoils in Transitional and Turbulent Regimes for Subsonic and Transonic Flows

### v1

@@ 5 Acknowledgement
This work used computational resources provided by ANVIL at Purdue University through allocation PHY240112 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, which is supported by U.S. National Science Foundation grants #2138259, #2138286, #2138307, #2137603, and #2138296 (Boerner et al. (2023)).
The first author would like to thank Dr. Shaowu Pan and Dr. Sicheng He for acquiring the NSF ACCESS allocation PHY240112: Scalable Scientific Machine Learning for Complex Physical Systems, which provided the computational resources on Anvil at Purdue University for the dataset generation.
@@ F.3 Prediction
We can determine how accurate our model is at reproducing the airfoil data field by calculating the $L_{2}$ norm error between the ground truth and predicted output. We can see two examples of higher $L_{2}$ norm in the $3^{\text{rd}}$ and $4^{\text{th}}$ airfoil in Figure 12, with error of 0.131 each.
Another way to examine the prediction quality of the model is to plot the absolute error $|C^{(i)}_{p,\text{real}}-C^{(i)}_{p,\text{pred}}|$ between the ground truth and the predicted output. This allows us to visualize areas of high error such as the shock present above the $3^{\text{rd}}$ and $4^{\text{th}}$ airfoil shown in  Figure 12. Often the model fails to fully represent the high pressure deltas of a shock resulting in high error present along the surface of a normal shock.
Despite this the model is capable of generating pressure coefficient data that can be used to locate the shock if not fully map it. This method also allows us to see areas of consistent error such as the field around the $1^{\text{st}}$ and $5^{\text{th}}$ airfoils.
Unlike the sharp discontinuities that appear near shocks this error is far more gradual and represents a failure of the model to determine what the average pressure coefficient values are far from the surface of the airfoil. It is possible that this is because the highest resolution input data makes up a small portion of the shown image. Since the unwrapping method extracts the pressure coefficient values as seen on the O-Grid, the model has a high bias to predicting values correctly closest to the airfoil surface.
As the distance to the surface of the airfoil increases, the resolution of the input data and thus the expected accuracy of the model decreases. Another possible cause of the high $L_{2}$ error in these figures is a low $L_{2}$ norm value for the real $C_{p}$. Since this term is in the denominator of the $L_{2}$ error formula, an extremely low value here will cause any small deviations in the numerator to result in high values for the error.

### v2

@@ 5 Acknowledgement
This work used resources provided by the NSF grants PHY240112: Scalable Scientific Machine Learning for Complex Physical Systems and PHY240208: Data-driven Laminar–turbulent Transistion Model Discovery and Design Optimization, which provided the computational resources on Anvil at Purdue University for the dataset generation.
@@ F.3 Prediction
We evaluate the model’s accuracy in reproducing the pressure coefficient ($C_{p}$) fields of airfoils by computing the $L_{2}$ norm of the error between the predicted and ground truth values. Notably, higher $L_{2}$ norm errors are observed for the 3rd and 4th airfoils in Figure 12, each with an error of approximately 0.131.
To further examine the model’s performance, we plot the absolute error field (Figure 12),
$\left|C^{(i)}_{p,\text{real}}-C^{(i)}_{p,\text{pred}}\right|,$ (16)
which allows visualization of localized discrepancies. As shown in Figure 12, regions of high error are found in the vicinity of shock waves, particularly over the 3rd and 4th airfoils. These errors typically arise because the model struggles to accurately represent the steep gradients in pressure associated with shocks, resulting in localized deviations along the shock front.
Despite this limitation, the model reliably captures the overall structure of the pressure field, including the location of shocks and discontinuities. This makes the output useful for shock detection, even if exact magnitudes are not fully recovered. In contrast to the abrupt errors near shocks, more diffuse and persistent error patterns are evident in the background fields, such as around the 1st and 5th airfoils. These are attributed to a broader modeling challenge: the failure to accurately reconstruct mean pressure values far from the airfoil surface.
One potential cause of this background error is the non-uniform resolution of the input data. The model receives unwrapped $C_{p}$ data from an O-grid representation, in which the regions closest to the airfoil surface are more densely sampled. As a result, the model becomes biased toward higher accuracy near the surface, while performance degrades with increasing distance.
Another contributing factor to elevated $L_{2}$ norm errors may be the small magnitude of the ground truth $C_{p}$ field in certain cases. Since the $L_{2}$ error is a relative measure, small denominator values in the error formula can amplify minor discrepancies in the numerator, leading to seemingly high error values.
To improve model performance and ensure consistency across varying geometries, two normalization strategies were applied. First, both the model outputs and the dataset were normalized to the range $[-1,1]$. Second, we recorded the minimum and maximum $C_{p}$ values from each simulation and appended them to the model’s latent space. This enabled the network to learn and predict the appropriate scaling factors directly from the geometry.
However, this approach introduces sensitivity: minor inaccuracies in the predicted extrema can propagate as large-scale errors across the entire field. This highlights a key limitation—without directly preserving the true range of the pressure coefficient values, the reconstructed fields may systematically deviate from the physical solution. Future work should explore alternative strategies to better preserve the physical fidelity of predicted $C_{p}$ distributions, particularly in the presence of shocks and weakly varying background fields.